OpenStudy (anonymous):

Algebra help please. simplify (2^5a^9b^5)(2a_4b^3)/2^3a^6

OpenStudy (babalooo):

\[\frac{ 2^5a^9b^5 \times 2a^4b^3}{ 2^3a^6 }\]

OpenStudy (anonymous):

Yes

OpenStudy (babalooo):

Well first of all, we can deal with the top first. We can combine those 2 terms since the base is the same. When multiplying with same base, just add exponents

OpenStudy (babalooo):

\[x^{a}\times x^{b} = x^{a+b}\]

OpenStudy (anonymous):

\[(2^5a^9b^5)(2a^-4b^3)/2^3a^6\]

OpenStudy (babalooo):

What are you suggesting?

OpenStudy (babalooo):

is my equation above wrong?

OpenStudy (babalooo):

is that a subtraction or...

OpenStudy (anonymous):

no...the numerator\[2a^-4 \not 2a^4\]

OpenStudy (anonymous):

\[is this correct? 2^5a^9b^5\times 2a^-4b^3 = 2^3a^6\]

OpenStudy (babalooo):

wait was my equation wrong?

OpenStudy (babalooo):

i don't get your 2a-4b^3 part

OpenStudy (anonymous):

The question is simplify \[(2^5a^9b^5)(2a^-4b^3)/2^3a^6\]

OpenStudy (babalooo):

\[\frac{ 2^5a^9b^5 \times 2a^{-4}b^3 }{ 2^3a^6 }\] Is this it?

OpenStudy (anonymous):

Yes

OpenStudy (babalooo):

and you got \[2^3a^6?\]

OpenStudy (anonymous):

No, not sure how to solve the problem

OpenStudy (babalooo):

do you get the rule i stated above?

OpenStudy (babalooo):

\[x^{a}\times x^{b} = x^{a+b}\]

OpenStudy (anonymous):

Which is a and b?

OpenStudy (babalooo):

example: (x^2)(x^3) = x^5 (xx)(xxx) = (xxxxx)

OpenStudy (anonymous):

ok

OpenStudy (babalooo):

that is just a general formula. let's start with the numerator. with the base 2

OpenStudy (babalooo):

in bracket one we have 2^5 and bracket 2^1 (2^5)*(2^1) = ?

OpenStudy (anonymous):

\[2^5a^9b^5\times2a^-4b^3\]

OpenStudy (babalooo):

yes? we are doing the base 2 first. essentially the top is all multiplication \[2^{5}2a^9a^{-4}b^5b^3\] is the same thing

OpenStudy (anonymous):

ok

OpenStudy (babalooo):

so what do you get for combining the base 2 values ?

OpenStudy (babalooo):

Or if you can do the whole numerator, that'd be great. I think it's pretty stright forward with the general rule and example I showed you

OpenStudy (anonymous):

\[2a^9+-4b^5+3\]

OpenStudy (babalooo):

what are you doing. can you explain?

OpenStudy (anonymous):

\[2a^5b^8\]

OpenStudy (babalooo):

CORRECT

OpenStudy (anonymous):

\[2a^5b^8/2^3a^6\]

OpenStudy (babalooo):

As for the denominator, it's subtraction of exponents

OpenStudy (babalooo):

\[\frac{ x^a }{ x^b } = x^{a-b}\]

OpenStudy (babalooo):

any exponents in the denominator brought to the numerator becomes negative

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[a^5-a^6= -1\]

OpenStudy (babalooo):

\[\frac{ x^a }{ x^b } = x^ax^{-b} = x^{a-b}\]

OpenStudy (babalooo):

just the exponents. what happen to your 2?

OpenStudy (babalooo):

\[a^{5-6} = a^{-1}\]

OpenStudy (babalooo):

don't forget your b

OpenStudy (anonymous):

\[2^2a^-1b^2\]

OpenStudy (babalooo):

are you sure about your b term?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry

OpenStudy (babalooo):

why is it to the power 2

OpenStudy (anonymous):

b^8

OpenStudy (babalooo):

also, i just noticed that your 2 has wrong power from your previous numerator solution

OpenStudy (babalooo):

you got 2a^5b^8 before. why is your 2 only to power 1?

OpenStudy (anonymous):

no..2^2a^-1b^8

OpenStudy (babalooo):

no i mean for the numerator only part

OpenStudy (anonymous):

yes

OpenStudy (babalooo):

\[2^{5} \times 2^1 = ?\]

OpenStudy (anonymous):

\[2^5+2^3\]

OpenStudy (babalooo):

no just the exponents

OpenStudy (anonymous):

how am confused now

OpenStudy (babalooo):

okay. your a and b are correct . it is the final answer. the a and b part

OpenStudy (babalooo):

but your 2 is part is incorrect.

OpenStudy (babalooo):

\[2^5a^9b^5 \times 2a^{-4}b^3\]

OpenStudy (anonymous):

remember the two b's in the numerator is b^5 and b^3

OpenStudy (babalooo):

that part is correct

OpenStudy (babalooo):

what about the 2 coefficient

OpenStudy (babalooo):

\[2^5 \times 2^1 = ?\] \[a^9 \times a^{-4} = a^5\] \[b^5 \times b^3 = b^8\]

OpenStudy (babalooo):

we're going back to just the numerator multiplication only

OpenStudy (babalooo):

no denominator yet. you solved the a and b correctly. we just need the 2 coefficient.

OpenStudy (anonymous):

ok

OpenStudy (babalooo):

You get it?

OpenStudy (anonymous):

\[2^5a^9+-4b^5+3= 2^5a^5b^8\]

OpenStudy (babalooo):

isn't it \[2^5a^9b^5 \times 2a^{-4}b^3\]

OpenStudy (anonymous):

ooh oh..you did subtraction with a?

OpenStudy (babalooo):

isn't that the equation?

OpenStudy (anonymous):

yes, the original equation

OpenStudy (babalooo):

exactly. so how did you get your 2 to the power of 5? for the numerator only?

OpenStudy (babalooo):

when 2 is alone, it is 2^1. 2^0 =1, 2^1=2, 2^2=4

OpenStudy (anonymous):

2 is't alone, it is 2^5

OpenStudy (babalooo):

you solved for a and b for the numerator and got \[a^5b^8\]

OpenStudy (babalooo):

what about the second part isn't it \[2a^{-4}b^3?\]

OpenStudy (anonymous):

yes, the second part is alone but the first isn't

OpenStudy (babalooo):

yes so it is \[2^5 \times 2^1= 2^?\]

OpenStudy (anonymous):

2^6

OpenStudy (babalooo):

imagine 2=c \[c^5 \times c^1 =?\]

OpenStudy (babalooo):

YAY! now do the denominator part to it as well! What do you get as your final answer?

OpenStudy (anonymous):

\[2^3a^-1b^8\]

OpenStudy (babalooo):

YES! That's the final answer.

OpenStudy (babalooo):

\[2^3a^{-1}b^3\] Once you get the hang of it. It isn't too bad right.

OpenStudy (anonymous):

yes, and thanks for your patience

OpenStudy (babalooo):

no worries. you can also expand the 2 \[8a^{-1}b^8\]

OpenStudy (babalooo):

2x2x2 = 8 if your teacher prefers that way

OpenStudy (anonymous):

ok